OpenStudy (anonymous):

what is the point-slope form that has a slope of 2 and passes through (1,-2)

OpenStudy (anonymous):

you just have to substitute the values to the formula: y-y1 = m (x-x1) so you will have: y+2 = 2(x-1) by solving: y+2 = 2x - 2 4 = 2x - y in standard form 0 = 2x - y in general form y = 2x - 4 in point slope form